Circular No. 7200

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 L3 (LINEAR) An apparently asteroidal fast-moving LINEAR object, posted on The

NEO Confirmation Page, was noted by F. B. Zoltowski, Woomera, S. Australia as having a tail ∼ 30′′ long in p.a. 100^{o} and a rather dense coma on June 13 and 14 CCD images. P. R. Holvorcem, Valinhos, Brazil, reports a ∼ 10′′ coma on his June 12 images. The discovery observation is given below, and additional astrometry appears on MPEC 1999-L34.

1999 UT  α_{2000}  δ_{2000}  Mag.
June 9.17148  13^{h}07^{m}18.^{s}98  −26^{o}31′47′′.8  18.3

Preliminary parabolic orbital elements by B. G. Marsden:

T = 1999 Dec. 31.319 TT  ω = 348.^{o}975
Ω = 140.019 }2000.0
q = 2.09910 AU  i = 166.070
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 11 13^{h}03.^{m}66 −25^{o}56.^{′}2 2.451 3.126 123.^{o}3 15.^{o}8 16.9
16 12 54.49 −24 20.9 2.497 3.086 116.4 17.1 16.9
21 12 46.47 −22 49.7 2.551 3.047 109.7 18.3 16.9
26 12 39.55 −21 23.9 2.610 3.009 103.1 19.2 16.9
July 1 12 33.64 −20 04.1 2.673 2.970  96.8 19.9 16.9

COMET C/1999 L2 (LINEAR) Additional astrometric observations are given on MPEC 1999-L33,

which also contains the following preliminary parabolic orbital elements:

T = 1999 Aug. 5.335 TT  ω = 179.^{o}055
Ω = 94.302 }2000.0
q = 1.90018 AU  i = 43.911
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 11 16^{h}35.^{m}40 +15^{o}33.^{′}0 1.133 2.021 140.^{o}2 18.^{o}8 16.3
16 16 29.54 +12 57.2 1.111 2.000 140.3 18.9 16.2
21 16 23.95 +10 04.1 1.096 1.982 139.5 19.4 16.2
26 16 18.83 + 6 56.5 1.088 1.965 138.1 20.2 16.1
July 1 16 14.34 + 3 37.8 1.088 1.950 135.9 21.3 16.1
1999 June 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green